# Strategies for the alignment of electronic states in quantum-dot tunnel-injection lasers and their influence on the emission dynamics

### v1

@@ Abstract
In quantum-dot tunnel-injection lasers, the excited charge carriers are efficiently captured from the bulk states via an injector quantum well and then transferred into the quantum dots via a tunnel barrier. The alignment of the electronic levels is crucial for the high efficiency of these processes and especially for the fast modulation dynamics of these lasers. In particular, the quantum mechanical nature of the tunneling process must be taken into account in the transition from two-dimensional quantum well states to zero-dimensional quantum dot states. This results in hybrid states, from which the scattering into the quantum-dot ground states takes place. We combine electronic state calculations of the tunnel-injection structures with many-body calculations of the scattering processes and insert this into a complete laser simulator. This allows us to study the influence of the level alignment and limitations due to inhomogeneous quantum-dot distributions. We find that the optimal alignment deviates from a simple picture in which the of the quantum-dot ground state energies are one LO-phonon energy below the injector quantum well ground state.
I Introduction
@@
Tunnel-injection (TI) lasers with quantum dots (QD) as active material promise improved emission properties, in particular faster modulation speed in optical data transmission networks. Sun et al. (1993); Bhattacharya and Ghosh (2002); Han and Asryan (2008); Bhowmick et al. (2014); Gready and Eisenstein (2010) In conventional QD lasers with current injection pumping, the pump process generates excited carriers in the bulk barrier states. Capturing these carriers into the QD states requires substantial energy dissipation via carrier-phonon-scattering in combination with carrier-carrier scattering. This limits the generation efficiency of excited carriers in the QD states and slows down the modulation speed of these lasers. The TI-laser design utilizes an injector quantum well (QW) that is directly coupled to the QDs via a tunnel barrier. The injector QW provides a much larger scattering cross section to the excited carriers driven by the carrier injection, thereby enhancing the capture rate of carrier from the bulk states. Tunnel coupling facilitates efficient transfer of excited carriers from the injector QW into the QD states.
@@
The design of TI-QD lasers requires proper alignment of the electronic states between the injector QW and the QDs to exploit the advantages of this system. The dominant carrier scattering process, which contributes to the tunnel-assisted carrier transfer from the injector QW to the QDs has been identified as carrier interaction with LO phonons. Sun et al. (1993); Bhattacharya and Ghosh (2002); Bhowmick et al. (2014); Gready and Eisenstein (2010); Michael et al. (2018); Lorke et al. (2018, 2022) Hence the original design goal was to engineer the structure in a way the QD ground states are positioned one LO-phonon energy below the band edge of the injector QW.
@@
A closer look reveals two challenges. Due to an inhomogeneous distribution of QD confinement energies, only some of the QDs have optimal tunnel coupling. Moreover, the tunnel coupling does not occur from the bottom of the injector QW. In Refs. Michael et al. (2018); Lorke et al. (2018, 2022) it was shown that the quantum mechanical coupling of the 2d and 0d subsystems has to be taken into account. It leads to hybrid states which are partly localized in the in the injector quantum well as well as in the QDs. The more elaborate picture for the transfer of excited carriers from the QW into the QD thus involves carrier scattering processes from such a hybrid state into the QD ground state through emission of LO phonons.
@@
We have developed microscopic models for the description of the excited carrier transfer from the injector QW to an inhomogeneously distributed QD ensemble. Starting from electronic state calculations for the coupled 2d-0d system, interaction matrix elements for the carrier-phonon interaction are determined. These are included in a quantum kinetic calculation of scattering rates, which account for polaron energy renormalizations and deviations from a strict energy conservation in terms of free-particle states for the scattering processes. Seebeck et al. (2005); Michael et al. (2018); Lorke et al. (2018, 2022) The results determine capture rates of excited electrons into the QDs as function of their detuning and the resulting carrier dynamics.
@@
In this paper we integrate our model for the carrier kinetics in tunnel-coupled injector QW and QD system into a complete laser simulator. The latter includes a spatially resolved description of the carrier transport in a structure with 6 injector QWs and the coupling of the excited carrier recombination in the QDs via spontaneous and stimulated emission to a self-consistent laser model.
II Theory
We consider a QD ensemble with size and composition fluctuations leading to a Gaussian distribution of ground state (s-shell) energies with a FWHM of 15meV. The maximum of the distribution is centered at an interband transition energy of 800meV (1550nm). The situation is sketched in Fig. 1 with the blue line. Fig. 1 shows three examples of the
@@ II Theory
QDs in the inhomogenoues ensemble with their ground state energy (solid colored line) and the associated hybridized injector QW states (colored shaded areas). These are no longer zero-dimensional states (associated with the former p-shell) Lorke et al. (2022), but have a two-dimensional component in the injector QW due to the tunnel coupling. Due to the differences in energetic alignment between QDs in the inhomogeneous ensemble, they couple differently to the injector QW.
In this work, we compare the current TI design, where the energetic alignment between the QD ground state and the injector QW ground state is tuned to the LO phonon energy (Alignment A in what follows), with a situation that we propose in this work, where the injector QW is shifted to lower energies in an effort to optimize the scattering efficiency (Alignment B).
To study switch-on and modulation properties of lasers based on TI structures as active material, we set up a model that includes both the microscopic carrier dynamics in the TI structure on the one hand, as well as the carrier transport though the bulk material on the other.
The carrier transport in the bulk material is described by
$\frac{d}{dt}N_{\text{res}}(z,t)|_{\text{transport}}=\frac{d}{dz}\left[D_{N}\frac{d}{dz}N_{\text{res}}(z,t)+\mu_{N}N_{\text{res}}E(z,t)\right]$ (1)
@@
In this equation $N_{res}$ is the three-dimensional carrier density of the bulk material, $D_{N}$ and $\mu_{N}$ are the diffusion coefficient and mobility of the carriers and $E(z,t)$ is the internal field obtained by a solution of Poisson’s equation for the carrier density.
The carriers in the bulk material are couples to the active region, consisting of 6 QW-QD layers that couple to the bulk carrier density
$\begin{split}\frac{d}{dt}N_{\text{res}}(z,t)&=\frac{d}{dt}N_{\text{res}}(z,t)|_{\text{transport}}-\frac{N_{\text{res}}(z,t)}{\tau_{res}}\\ &-N_{\text{res}}(z,t)\sum_{i=1}^{M}\frac{N_{QD}^{i}}{\tau^{i}_{cap}}(1-f^{i;e,h})\\ &+\sum_{i=1}^{M}\frac{N_{QD}^{i}}{\tau^{i}_{\text{esc}}}f^{i;e,h}\left(1-\frac{N_{\text{res}}}{D_{res}}\right)\\ &-\sum_{j}\frac{N_{\text{res}}}{\tau_{cap,IW}}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\\ &+\sum_{j}\frac{N^{j}_{IW}}{\tau_{esc,IW}}\left(1-\frac{N_{\text{res}}}{D_{res}}\right)\leavevmode\nobreak\ .\end{split}$ (3)
Here, $D_{res}$ ($D_{IW}$) is the density of states of the bulk (injector QW), $N^{j}_{IW}$ the injector QW carrier density of layer $j$, $f^{i;e,h}$ the electron and hole occupation of the of the QD s-shell in the i-th sub-ensemble. $\tau^{i}_{cap,esc}$ is the capture (escape) scattering time from the bulk into the QD and $\tau_{cap,esc,IW}$ is the capture (escape) scattering time between bulk and injector QW.
@@
$\begin{split}\frac{d}{dt}N^{j}_{IW}&=\frac{N^{j}_{IW}}{\tau_{IW}}-\frac{N^{j}_{IW}}{\tau_{\text{tun}}}\sum_{i=1}^{M}N_{QD}^{i}(1-f^{i;e,h})\\ &+\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\sum_{i=1}^{M}2N_{QD}^{i}\frac{f^{i;e,h}}{\tau_{tun,b}}\\ &+\frac{N_{\text{res}}(z_{W},t)}{\tau_{cap.IW}}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\\ &-\frac{N^{j}_{IW}}{\tau_{esc,IW}}\left(1-\frac{N_{\text{res}}(z_{W},t)}{D_{res}}\right)\leavevmode\nobreak\ ,\end{split}$ (4)
with $\tau_{tun}$ is the tunnel scattering time between the hybridized injector QW state and the QD s-shell, as describe in our previous publication Lorke et al. (2022).
The QD populations $f^{i;e,h}$ follow from the semiconductor Bloch equations Haug and Koch (2004)
@@
$\begin{split}\frac{d}{dt}f^{i;e,h}=&-\gamma_{c}f^{i;e,h}+\frac{N_{\text{res}}(z_{D},t)}{2\tau_{c}ap}(1-f^{i;e,h})\\ &-\frac{f^{i;e,h}}{\tau^{i}_{\text{esc}}}(1-\frac{N_{\text{res}}(z_{D},t)}{D_{res}})\\ &-2\text{Im}\left(\Omega_{\text{lase}}\psi^{i,\ast}\right)+\sum\limits_{j}\frac{N^{j}_{IW}}{2\tau_{\text{tun}}}(1-f^{i;e,h})\\ &-\sum\limits_{j}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\frac{f^{i;e,h}}{\tau_{tun,b}}\end{split}$ (5)
with the interband polarization
@@
$\frac{d}{dt}\psi^{i}=-(i\omega+\gamma_{h})\psi^{i}-i\frac{\Omega_{\text{lase}}}{\hbar}\left(1-f^{i;e}-f^{i;h}\right)\leavevmode\nobreak\ ,$ (6)
where $\gamma_{h}$ is the homogeneous linewidth.
III Results
@@
We consider an inhomogeneous QD ensemble with size and composition fluctuations, leading to a Gaussian distribution of ground state (s-shell) energies. The QD layers are separated from the injector QW by a thin 2.1 nm thick tunnel barrier. The active material of the laser consists of 6 QW-QD layers separated by 10.5 nm wide spacers layers.
Figure 2: Inverse tunnel scattering time as a fucntion of QD s-shell energy for the present experimental alignment (red line) to the QD distribution (black line) as well as the optimized alignment (blue line).
Previously Lorke et al. (2022), we have demonstrated that current TI designs (alignment A) couple rather inefficiently to the injector QW. This is shown (red line) in Fig. 2 together with the QD distribution (black line). Here, the dashed line refers to the situation, where the p-shell is hybridized, while for the dotted line the p-shell is confined and the d-shell hybridizes with the QW. For details, see Ref. Lorke et al. (2022). In comparison our proposed alignment B (blue line in Fig. 2) has an improved carrier scattering at the maximum of the QD distribution. Such an alignment could be generated by, e.g., modifying the injector QW width or concentration profile. One of the main reasons for the improved scattering rates in alignment B is that the most efficient carrier scattering does not occur from the bottom of the QW but from higher lying states due to symmetry considerations. Hence, even in a independent particle picture the QW needs to be tuned accordingly.
@@
In Fig. 3 input-output characteristics are shown for alignment A and B. Notably, the proposed alignment B has a higher threshold current than the current alignment A. The reason is that carriers are more evenly distributed over the 6 active layers with the optimized alignment (see Fig. 4), leading to a higher carrier density to be accumulated in the device before lasing occurs. This is caused by the additional carriers in the injector QWs of the outer layers, that need to be populated for carrier scattering into the QD ground states to occur.
@@
Figure 3: Input/Output characteristics for TI laser devices with 6 active layers and current (red) and optimized (blue) alignment between the injector QW and the QD ensemble
@@
To investigate this further, the total carrier density $(N_{res}+N_{IW}+N_{QD}f^{i;e})$ is shown in Fig. 4 for a pump power 10% above the laser threshold for both devices. Electrons (Holes) are generated by the contact in this Figure at the left (right) edge of the device. For the device with the current experimental alignment A, the slower tunnel scattering rate leads carriers more localized in the central layers of the device.
Figure 4: total electron density as a function of position and time for the current (blue) and optimized (red) alignment situation.
@@
While the higher threshold seems to indicate that the proposed "optimized" alignment B performs worse, the situation is reversed when investigating dynamical laser properties such as the switch-on process. The dynamic of the output power is shown in Fig. 5. For better visibility, only the first 300 ps are shown, whereas the simulation is run to reach steady state.
@@
Figure 5: Output power as a function of time for a pump power 10% above the laser threshold for the current alignment (red) and the optimized alignment (blue)
@@
As the carrier spread out more evenly over the 6 layer structure, the onset of lasing shows a small delay of about 10ps Clearly, however, the optimized structure shows a significantly faster switch-on process. Comparing the results to laser rate equation models Coldren and Corzine (1995) reveals that the optimized structure should be close to critical damping of the relaxation oscillations Lorke et al. (2011), as only half an oscillation cycle is visible. This finding already leads us to expect enhanced modulation capabilities of the optimized structures in comparison with current experimentally realized ones.
IV Acknowledgment

### v3

@@ Abstract
In quantum-dot tunnel-injection lasers, the excited charge carriers are efficiently captured from the bulk states via an injector quantum well and then transferred into the quantum dots via a tunnel barrier. The alignment of the electronic levels is crucial for the high efficiency of these processes and especially for the fast modulation dynamics of these lasers. In particular, the quantum mechanical nature of the tunneling process must be taken into account in the transition from two-dimensional quantum well states to zero-dimensional quantum dot states. This results in hybrid states, from which the scattering into the quantum-dot ground states takes place. We combine electronic state calculations of the tunnel-injection structures with many-body calculations of the scattering processes and insert this into a complete laser simulator. This allows us to study the influence of the structural design and the resulting electronic states as well as limitations due to inhomogeneous quantum-dot distributions. We find that the optimal electronic state alignment deviates from a simple picture in which the of the quantum-dot ground state energies are one LO-phonon energy below the injector quantum well ground state.
@@
Tunnel-injection (TI) lasers with quantum dots (QD) as an active material promise improved emission properties, in particular faster modulation speed in optical data transmission networks. Sun et al. (1993); Bhattacharya and Ghosh (2002); Han and Asryan (2008); Bhowmick et al. (2014); Gready and Eisenstein (2010) In conventional QD lasers with current injection pumping, the pump process generates excited carriers in the bulk barrier states. Capturing these carriers into the QD states requires substantial energy dissipation via carrier-phonon-scattering in combination with carrier-carrier scattering. This limits the excitation efficiency for electrons and holes in the QD states and slows down the modulation speed of these lasers. The TI-laser design utilizes an injector quantum well (QW) that is directly coupled to the QDs via a tunnel barrier. The injector QW provides a much larger scattering cross section to the excited carriers driven by the carrier injection, thereby enhancing the capture rate of carrier from the bulk states. Tunnel coupling facilitates efficient transfer of excited carriers from the injector QW into the QD states.
@@
The design of TI-QD lasers requires proper alignment of the electronic states between the injector QW and the QDs in order to exploit the advantages of this system. The dominant carrier scattering process, which contributes to the tunnel-assisted carrier transfer from the injector QW to the QDs has been identified as carrier interaction with LO phonons. Sun et al. (1993); Bhattacharya and Ghosh (2002); Bhowmick et al. (2014); Gready and Eisenstein (2010); Michael et al. (2018); Lorke et al. (2018, 2022) Hence the original design goal was to engineer the structure in a way that the QD ground states are positioned one LO-phonon energy below the band edge of the injector QW.
@@
A closer look reveals two challenges. Due to an inhomogeneous distribution of QD confinement energies, only some of the QDs have optimal tunnel coupling. In the presently used material systems like InGaAs, the inhomogeneous broadening of energies has a FWHM of 15 to 25 meV in comparison to the LO-phonon energy of 36 meV. In addition, tunnel coupling into the QDs is inefficient for the lowest conduction band states of the injector QW. Optimal tunnel coupling for the conduction-band states is linked to hybrid states between the first excited QD states and injector QW states with finite carrier momentum. The last two statements have been derived from a quantum mechanical treatment of the electronic coupling between the 2d and 0d single-particle states, Michael et al. (2018); Lorke et al. (2018, 2022) which leads to hybrid states that are partly localized in both the injector QW as well as in the QDs. The transfer of excited charge carriers from the QW into the QD then takes place in the form of scattering processes between these hybrid states and the QD ground state.
@@
We have developed a microscopic description for the excited charge carrier transfer from the injector QW to an inhomogeneously distributed QD ensemble. Calculations of the electronic states in the coupled 2d-0d system serve as a starting point. These provide single-particle states and interaction matrix elements for the carrier-phonon interaction and carrier-carrier Coulomb interaction. While Boltzmann-type scattering rates in kinetic equations could have been used as a subsequent step, we find that a quantum kinetic calculation of scattering rates provide a more appropriate description, as it accounts for polaron energy renormalizations and deviations from a strict energy conservation in terms of free-particle states. Seebeck et al. (2005); Michael et al. (2018); Lorke et al. (2018, 2022) The results determine capture rates of excited electrons into the QDs as function of their detuning and the resulting carrier dynamics. Here the carrier scattering through emission of LO phonons has been identified as the dominant process while carrier-carrier Coulomb scattering additionally contributes.
@@
In this paper we integrate our description of the carrier kinetics in tunnel-coupled injector QW and QD system into a complete laser simulator. The latter includes a spatially resolved modeling of the carrier transport in a structure with 6 injector QWs and the coupling of the excited carrier recombination in the QDs via spontaneous and stimulated emission to a self-consistent laser model. We consider a QD ensemble with size and composition fluctuations leading to a Gaussian distribution of ground state (s-shell) energies with a FWHM of 15meV. The maximum of the distribution is centered at an interband transition energy of 800meV (1550nm). The situation is sketched in Fig. 1 with the blue line.
@@
In addition three examples for QDs in the inhomogenoues ensemble are shown with their ground state energy (solid colored line) and the associated hybridized injector QW-QD states (colored shaded areas). The latter are the new eigenstates due to tunnel coupling, which are formed from zero-dimensional excited QD states and two-dimensional injector QW states. For a given QD, the hybridization takes place for a range of QW energies. The energy with strongest coupling is lowered in comparison to the uncoupled excited QD state, as the tunnel coupling weakens the confinement of this state. The hybridized states are calculated for the individual QDs and vary to some extent within the inhomogeneous ensemble. This applies in particular to the wave function symmetry when involving low carrier momentum injector QW states. Lorke et al. (2022)
In this work, we compare the previously suggested TI design, in which the energetic distance between the QD ground state and the injector QW ground state is tuned to the LO phonon energy (alignment A in what follows), with a situation that we propose in this work, where the injector QW is shifted to lower energies in an effort to optimize the scattering efficiency (alignment B). To study switch-on and modulation properties of lasers with TI structures as active material, we use a model that includes both the microscopic carrier dynamics in the TI structure, as well as the carrier transport through the bulk material.
The charge carrier transport in the bulk material regions, including the injection of carriers and their distribution through the structure via drift and diffusion is described on a classical level by
$\begin{split}\frac{d}{dt}N_{\text{res}}&(z,t)|_{\text{transport}}\\ &=\frac{d}{dz}\left[D_{N}\frac{d}{dz}N_{\text{res}}(z,t)+\mu_{N}N_{\text{res}}E(z,t)\right]\end{split}$ (1)
@@
In this equation $N_{\text{res}}$ is the three-dimensional carrier density of the bulk material, $D_{N}$ and $\mu_{N}$ are the diffusion coefficient and mobility of the carriers and $E(z,t)$ is the internal field obtained from a solution of Poisson’s equation for the carrier density.
The net result for the carrier dynamic in the bulk material contains coupling to the active region, as well as a loss term for the reservoir due to spontaneous recombination with the time constant $\tau_{\text{res}}$ ,
$\begin{split}\frac{d}{dt}N_{\text{res}}(z,t)&=\frac{d}{dt}N_{\text{res}}(z,t)|_{\text{transport}}-\frac{N_{\text{res}}(z,t)}{\tau_{\text{res}}}\\ &-N_{\text{res}}(z,t)\sum_{i=1}^{M}\frac{N_{\text{QD}}^{i}}{\tau^{i}_{\text{cap}}}(1-f^{i;e,h})\\ &+\sum_{i=1}^{M}\frac{N_{QD}^{i}}{\tau^{i}_{\text{esc}}}f^{i;e,h}\left(1-\frac{N_{\text{res}}}{D_{\text{res}}}\right)\\ &-\sum_{j}\frac{N_{\text{res}}}{\tau_{\text{cap,QW}}}\left(1-\frac{N^{j}_{\text{QW}}}{D_{\text{QW}}}\right)\\ &+\sum_{j}\frac{N^{j}_{\text{QW}}}{\tau_{\text{esc,QW}}}\left(1-\frac{N_{\text{res}}}{D_{\text{res}}}\right)\leavevmode\nobreak\ .\end{split}$ (3)
The second and third lines of Eq. (3) describe direct scattering of the carriers from/to the bulk to/from the QD states with times constants $\tau^{i}_{\text{cap}}$ and $\tau^{i}_{\text{esc}}$, respectively, where $f^{i;e,h}$ the electron and hole occupation of the of the QD s-shell. Originally, the sum over $i$ numbers the QDs with inhomogeneous distribution of transition energies. For practical purposes, we group QDs with similar electronic properties, where $M$ is the number of sub-ensembles and $N_{\text{QD}}^{i}$ is the number of QDs in the sub-ensemble.
The fourth and fifth rows of Eq. (3) describe capture and escape form the bulk states into the injector QW with carrier density $N^{j}_{\text{QW}}$ and density of states $D_{\text{QW}}$. The respective time constants are $\tau_{\text{cap,QW}}$ and $\tau_{\text{esc,QW}}$. The index j numbers the different injector QWs.
@@
$\begin{split}\frac{d}{dt}N^{j}_{\text{QW}}&=-\frac{N^{j}_{\text{QW}}}{\tau_{\text{QW}}}\\ &-\sum_{i=1}^{M}\frac{N^{j}_{\text{QW}}}{\tau^{i}_{\text{tun}}}N_{\text{QD}}^{i}(1-f^{i;e,h})\\ &+\sum_{i=1}^{M}\left(1-\frac{N^{j}_{\text{QW}}}{D_{\text{QW}}}\right)2N_{\text{QD}}^{i}\frac{f^{i;e,h}}{\tau^{i}_{\text{tun,b}}}\\ &+\frac{N_{\text{res}}(z^{j}_{\text{QW}},t)}{\tau_{\text{cap,QW}}}\left(1-\frac{N^{j}_{\text{QW}}}{D_{\text{QW}}}\right)\\ &-\frac{N^{j}_{\text{QW}}}{\tau_{\text{esc,QW}}}\left(1-\frac{N_{\text{res}}(z^{j}_{\text{QW}},t)}{D_{\text{res}}}\right)\leavevmode\nobreak\ ,\end{split}$ (4)
where the first line represents the loss of carriers due to spontaneous recombination with the time constant $\tau_{\text{QW}}$ The second and third line account for carrier scattering between the hybridized injector QW states and the QD s-shell with the associated rates $1/\tau^{i}_{\text{tun}}$ and $\tau^{i}_{\text{tun,b}}$, as describe in our previous publication Lorke et al. (2022) and in the Supplement. The fourth and fifth line are the counter terms to the respective contributions in Eq. (3). Here $z^{j}_{\text{QW}}$ is the position of the $j$-th injector QW.
The ground-state population of the $i$-th QD, $f^{i;e,h}$, is calculated from the semiconductor Bloch equations Haug and Koch (2004) with the added contributions due to the carrier scattering as introduced above. The QD population dynamics
@@
$\begin{split}\frac{d}{dt}f^{i;e,h}=&-\frac{f^{i;e}f^{i;h}}{\tau_{\text{sp}}}-2\text{Im}\left(\Omega_{\text{lase}}\psi^{i,\ast}\right)\\ &+\frac{N_{\text{res}}(z^{i}_{\text{QD}},t)}{2\tau^{i}_{\text{cap}}}(1-f^{i;e,h})\\ &-\frac{f^{i;e,h}}{\tau^{i}_{\text{esc}}}(1-\frac{N_{\text{res}}(z^{i}_{\text{QD}},t)}{D_{\text{res}}})\\ &+\sum\limits_{j}\frac{N^{j}_{\text{QW}}}{2\tau^{i}_{\text{tun}}}(1-f^{i;e,h})\\ &-\sum\limits_{j}\left(1-\frac{N^{j}_{\text{QW}}}{D_{\text{QW}}}\right)\frac{f^{i;e,h}}{\tau^{i}_{\text{tun,b}}}\end{split}$ (5)
is coupled to the optical interband polarization
@@
$\frac{d}{dt}\psi^{i}=-(i\omega+\gamma_{h})\psi^{i}-i\frac{\Omega_{\text{lase}}}{\hbar}\left(1-f^{i;e}-f^{i;h}\right)\leavevmode\nobreak\ .$ (6)
Here $\tau_{\text{sp}}$ is the spontaneous emission time and $\gamma_{h}$ is the homogeneous linewidth. Furthermore, $\Omega_{\text{lase}}=dE(z^{i}_{\text{QD}},t)$ contains the interband dipole coupling matrix element $d$ as well as the laser field $E$ at the QD position $z^{i}_{\text{QD}}$. The laser field $E(z,t)$ can be written in single-mode approximation as
$E(z,t)$ $=$ $\frac{1}{2}\mathscr{E}(t)u(z)e^{-i\omega_{c}t}+c.c.$ (7)
with
$\frac{d\mathscr{E}}{dt}$ $=$ $-\gamma_{c}\mathscr{E}+\frac{i\omega_{c}}{\varepsilon_{b}}\sum_{i=1}^{M}N_{\text{QD}}^{i}d^{*}\psi^{i}+\frac{f^{i;e}f^{i;h}}{\tau_{\text{sp}}}\leavevmode\nobreak\ ,$ (8)
where $u(z)$ is the mode function, $\omega_{c}$ is the laser frequency, and $\gamma_{c}$ describes the cavity field loss rate.
@@
We consider an inhomogeneous QD ensemble with size and composition fluctuations, described by a Gaussian distribution of ground state (s-shell) energies. The QD layers are separated from the injector QW by a 2.1 nm tunnel barrier. The active material of the laser consists of 6 QW-QD layers separated with 10.5 nm wide spacers layers.
Figure 2: Carrier capture rates from the injector QW into the QD as function of the QD ground-state energy relative to the QW zero-momentum state. Dashed lines correspond to the situation in which the p-shell is hybridized, while for the dotted line the p-shell is confined and the d-shell hybridizes with the QW. Red lines (alignment A, original design recommendation) refer to the bottom energy scale and blue lines (alignment B, new design recommendation) to the top energy scale. For both cases the inhomogeneous QD distribution is shown as black line.
Carrier scattering rates into the QD ground state are calculated by means of a quantum-kinetic theory which includes renormalizations of the electronic energies due to the electron-phonon coupling (polaronic effects Seebeck et al. (2005)). In this way we obtain an accurate description of carrier scattering by interaction with LO phonons, which according to our studies in Ref. Michael et al. (2018) forms the dominant scattering channel and is even more efficient than carrier-carrier Coulomb scattering, which we also considered there. We use three-dimensional quantum mechanical calculations of the electronic states in the coupled injector QW - tunnel barrier - quantum dot system using the nextnano3 software package nex . The resulting wavefunctions are directly applied in calculations of the interaction matrix elements for the scattering processes. On this basis, the influence of the TI design on the charge carrier dynamics in the laser is considered on a microscopic level. Further details can be found in the supplementary material. The resulting carrier capture rates into the QD ground state as function of its energy relative to the bottom of the injector QW conduction band are shown in Fig. 2 together with the QD distribution (black line). We find that the capture efficiency is small, if the QD s-shell energy is one LO-phonon energy below the bottom of the injector QW. This is a result of the wavefunction symmetry. The capture efficiency steadily increases for higher energies until it drops due to the dereasing carrier population of these states, providing a capture bandwidth of more than 15 to 20 meV. If the detuning of the QD ground state below the bottom of the injector QW is too large (the p-shell is below the band edge of the injector QW), the p-shell can no longer hybridize and the QD d-shell takes over the carrier capture.
The red lines in Fig. 2 represent the original design proposal for TI-QD lasers. When comparing the capture rate relative to the inhomogeneous QD distribution (black line), the majority of QDs does not benefit from efficient hybridization. For the high-energy tail of the QD distribution, the QD p-shell states are more than 15 meV above the injector QW band edge and hybridize increasingly efficient, thus providing higher capture rates. The low-energy tail of the QD distribution has d-shell energies $\gtrsim$ 15meV above the QW band edge, thus allowing for their efficient hybridization. Reflecting the tail of the carrier population in the injector QW, their capture rate increases for smaller energies.
Our main finding is that the scattering reaches its maximum if the QD ground states couple via LO phonons with 36 meV energy to injector QW states approximately 15 meV above their conduction band edge. This leads to our proposed alignment B (blue lines in Fig. 2), which places the injector QW energetically closer to the QD ensemble, thereby enabling improved carrier scattering at the maximum of the QD distribution. This alignment could be achieved, for example, by changing the injector QW width or concentration profile. We would like to point out that this effect is mainly the result of wavefunction engineering and can be also verified with a perturbative description of carrier scattering, based on Fermi’s golden rule.
@@
In Fig. 3 calculated input-output characteristics are shown for alignments A and B. It is worth noting that the proposed alignment B has a higher threshold current than the previously suggested alignment A. This is the result of a more uniform distribution of the carrier population across the 6 active layers with the optimized alignment B (see Fig. 4), which requires a higher carrier density in the device before lasing can occur. Specifically, additional carriers are need in the outer layers so that they can contribute better to laser emission.
@@
Figure 3: Input/Output characteristics for TI laser devices with 6 active layers using the previously recommended (red) and optimized (blue) alignment between the injector QW and the QD ensemble
@@
To investigate this further, the total carrier density $(N_{\text{res}}+N_{\text{QW}}+N_{\text{QD}}f^{i;e})$ is shown in Fig. 4 for a pump power 10% above the laser threshold for both devices. Electrons (holes) are generated by contacts at edge of the device corresponding to the left (right) side in this figure. For alignment A, the slower carrier capture rate leads to carriers being more localized in the central layers of the device.
Figure 4: Total electron density as a function of position for the two discussed alignments of the QD energy distributions relative to the injector QW.
@@
While the higher threshold seems to indicate that the proposed "optimized" alignment B performs worse, the situation is reversed when investigating dynamical laser properties such as the switch-on process. The time evolution of the output power is shown in Fig. 5. For better visibility, only the first 300 ps are displayed, whereas the simulation is carried out until the steady state is reached.
@@
Figure 5: Output power as function of time using a pump power 10% above the laser threshold for the two discussed alignments of the QD energy distributions relative to the injector QW.
@@
The realization of a spatially more spread out carrier distribution for alignment B is associated with a small delay of about 10ps in the onset of lasing. However, the optimized structure reaches the maximum intensity significantly faster, i.e., the switch-on process is improved. A comparison of the results to laser rate equation models Coldren and Corzine (1995) indicates that the optimized structure should be close to critical damping of the relaxation oscillations, Lorke et al. (2011) as only half an oscillation cycle is visible. This finding suggests an improved modulation capability of the optimized structures compared to those currently realized experimentally. We would like to point out the main difference stems not from the inclusion of z-dependent transport, but from a the modification of the energetic tuning between QDs and injector QW, that leads to more efficient carrier capture into the QD ground state. This effect results in a more efficient provision of excited charge carriers for the laser transition and hence to more efficient modulation properties.
In our calculations, we use a low inhomogeneous broadening of 15 meV, which corresponds to the currently best possible results. Higher inhomogeneous broadening reduces the number of QD that efficiently couple into the laser mode. For a non-optimal alignment, higher inhomogeneous broadening increases the fraction of QDs with better carrier capture. However, as these will be detuned from the laser mode positioned at the inhomogeneous distribution peak, the first effect mentioned still leads to a deterioration the laser emission properties.
In a recent paper, Khanonkin et al. (2022) a TI-QD laser with alignment A was compared to a non-TI laser with nominally identical QDs. The experimental results demonstrate lower threshold currents for the TI-QD laser but also significantly lower bandwidth compared to the non-TI QD laser. Although this is a different comparison, it corroborates that TI-QD lasers with alignment A have beneficial threshold but unfavorable modulation properties, as only parts of the QD ensemble experiences optimal coupling to the injector QW.
In conclusion, we have shown that the modulation performance of TI-based QD laser structures can be substantially improved by optimizing the energetic alignment between QD ensemble and injector QW. Faster carrier scattering rates result in a more efficient carrier transport into the lasing ground-state and shorten the switch-on time. Simultaneously this leads to a concentration of carriers in the central QD layers of the device, which impacts the switch-on current.
Supplementary Material
The supplementary material contains a detailed description of the theoretical methods used to calculate the carrier scattering rates.
I Acknowledgment